# Bias and Generalizability of Foundation Models across Datasets in Breast Mammography

### v1

@@ 2 Methodology
We investigate several strategies to promote domain invariance and mitigate biases. Specifically, we explored two categories of techniques to balance classifiers’ performance across datasets: i) domain-adaptation strategies (DANN, FairDisCO, and FADES) primarily designed to learn domain-invariant representations, and ii) fairness-aware strategies (GroupDRO, MOE) which explicitly focus on reducing performance disparities across subgroups. Table 1 summarizes the formulations of each method.
@@
Domain Adversarial Neural Network(DANN)
@@
Fair Disentanglement with Contrastive Learning (FairDisCO) employs adversarial and contrastive learning. It encourages samples from different domains with the same label to be close in a new feature space to mitigate bias [9].
Fair Disentanglement with Sensitive Relevance (FADES) penalizes features predictive of domain while maintaining those correlated with both domain and target tasks. It integrates total correlation (TC), conditional mutual information (CMI), and adversarial loss to minimize domain information leakage [16].
Group Distributionally Robust Optimization (GroupDRO) explicitly optimizes for the worst-case performance across domains. By re-weighting the loss based on each domain’s performance, GroupDRO ensures that the model does not favor majority groups at the expense of under-represented ones [28].
Mixture-of-Experts (MOE) uses a set of expert classifiers , each specializing in different domains, and combines their outputs through a gating mechanism . This allows the model to adaptively leverage domain-specific expertise while benefiting from a shared representation, as described in [19].
@@ 3 Experiments and Results
Images were preprocessed using the framework proposed by [11]. We used a rule-based approach to crop images according to the breast ROI. We set values less than 40 to 0 and eliminated consistently identical rows and columns, supposing these denote background. The final images had a size of 1,520×912. Experiments were implemented in Python v3.10 using Pytorch v2.4.1. Individual classifiers were trained on each dataset, and Unified on the aggregated datasets for two tasks: diagnosis and breast density. We searched for optimal batch size (8, 16, or 32) and learning rate (1e-3, 1e-4, or 1e-5) using 3-fold cross-validation within the training set. Hyperparameters giving the best accuracy after 20 epochs were then used for training on the whole set for 50 epochs. We used the same hyperparameter optimization for all mitigation strategies, except for FADES. Due to its computational cost, we fixed the batch size to 32, the learning rate to 1e-4, and trained for 30 epochs. Technical details specific to each strategy are reported in the code and will be publicly available upon acceptance.
@@ 3.2 Robustness of classifiers to domain-shift
Tab. 3 shows the performance of classifiers trained using features extracted from each FM for diagnosis and breast density classification. For both tasks, Unified classifiers trained on features from MammoCLIP outperform those based on GLORIA (, average relative improvement of +15.3%), CLIP (, +9.7%), and DinoV2 (, +13.3%), highlighting the advantages of pre-training on modality-specific data compared to domain-related (i.e. X-rays) or natural images. It is worth mentioning that MedCLIP-based classifiers exhibit notably poor performance across all tasks, suggesting that the extracted features may be predominantly noisy. Overall, classifiers have high performance when tested on the same dataset they were trained on (Indiv. (internal)), with average F1 scores of 0.73 and 0.53 for MammoCLIP on diagnosis and breast density, respectively. However, when tested on other datasets (Indiv. (external)), a drastic drop in F1 scores (up to for diagnosis) was remarkable, potentially due to overfitting of dataset-specific characteristics encoded in MammoCLIP’s features.
@@ 3.3 Effectiveness of bias and domain-adaptation strategies
Unified classifiers show performance similar to individual classifiers on their test sets (internal), with F1 score variations of ±15%. Aggregating datasets effectively improves generalization compared to individual classifiers (overall F1 score improved by and , with Unified vs. Indiv.). However, Unified classifiers exhibit performance disparities, with F1 score standard deviations of 0.15 across test datasets, suggesting that such aggregation cannot fully mitigate biases. DANN shows slightly lower overall F1 scores than Unified classifiers, especially for the diagnosis task. While this technique aims to learn domain-invariant representations, it seems to do so at the cost of overall performance. Similar observations can be made for FairDisCO, FADES, and MOE. For Breast density, GroupDRO produces consistently tighter F1 score distributions than Unified classifiers with standard deviations of 0.08, 0.06, and 0.08 for MammoCLIP, CLIP, and DinoV2, respectively, indicating reductions of disparities across test datasets. Additionally, it outperforms the Unified classifier across FMs, e.g. MammoCLIP with a relative improvement of 12% (). For diagnosis, mitigation strategies do not improve overall performance nor reduce disparities compared to Unified, likely due to variability in diagnostic label availability across datasets, e.g. no benign samples for MMD.
@@ 3.4 Bias and domain-adaptation in under-represented subgroups
Certain subgroups are under-represented in our datasets, e.g. breast density classes A and D representing 11 and 9% of the dataset and age <40 (10%) and >70 (10%), and are unequally represented across datasets (see Tab. 2). Fig.2 (top) illustrates prediction disparities across datasets and breast density classes, where indicates fair performance. GroupDRO and DANN achieve the most fair performance for all breast density classes (, and across breast density classes and datasets), aligning with DANN’s domain-invariant feature learning strategy. However, for DANN, this fairness comes at the cost of performance (see Tab. 3). GroupDRO stabilizes performance across breast density classes, notably improving prediction for class A, which other classifiers struggled with. This ability to learn across domains, while favoring under-represented subgroups, is critical for extreme breast densities (A, D) and age groups due to their strong interplay [8]. Fig. 2 (bottom) presents samples from different subgroups and classifiers’ successes and failures in breast density classification. Variations in contrast, texture, and patterns across classes and datasets may introduce spurious correlations, underscoring the need for fairness-aware strategies. GroupDRO seems an effective strategy to mitigate these biases and could be further refined to incorporate more fine-grained attributes.

### v2

@@ 2 Methodology
We investigate several strategies to promote domain invariance and mitigate biases. Specifically, we explored two categories of techniques to balance classifiers’ performance across datasets: i) domain-adaptation strategies (DANN
, FairDisCO
, and FADES
) primarily designed to learn domain-invariant representations, and ii) fairness-aware strategies (GroupDRO
, MOE
) which explicitly focus on reducing performance disparities across subgroups. Table 1 summarizes the formulations of each method.
@@
Domain Adversarial Neural Network(DANN
@@
Fair Disentanglement with Contrastive Learning (FairDisCO
) employs adversarial and contrastive learning. It encourages samples from different domains with the same label to be close in a new feature space to mitigate bias [9].
Fair Disentanglement with Sensitive Relevance (FADES
) penalizes features predictive of domain while maintaining those correlated with both domain and target tasks. It integrates total correlation (TC), conditional mutual information (CMI), and adversarial loss to minimize domain information leakage [16].
Group Distributionally Robust Optimization (GroupDRO
) explicitly optimizes for the worst-case performance across domains. By re-weighting the loss based on each domain’s performance, GroupDRO
ensures that the model does not favor majority groups at the expense of under-represented ones [28].
Mixture-of-Experts (MOE
) uses a set of expert classifiers , each specializing in different domains, and combines their outputs through a gating mechanism . This allows the model to adaptively leverage domain-specific expertise while benefiting from a shared representation, as described in [19].
@@ 3 Experiments and Results
Images were preprocessed using the framework proposed by [11]. We used a rule-based approach to crop images according to the breast ROI. We set values less than 40 to 0 and eliminated consistently identical rows and columns, supposing these denote background. The final images had a size of 1,520×912. Experiments were implemented in Python v3.10 using Pytorch v2.4.1. Individual classifiers were trained on each dataset, and Unified
on the aggregated datasets for two tasks: diagnosis and breast density. We searched for optimal batch size (8, 16, or 32) and learning rate (1e-3, 1e-4, or 1e-5) using 3-fold cross-validation within the training set. Hyperparameters giving the best accuracy after 20 epochs were then used for training on the whole set for 50 epochs. We used the same hyperparameter optimization for all mitigation strategies, except for FADES
. Due to its computational cost, we fixed the batch size to 32, the learning rate to 1e-4, and trained for 30 epochs. Technical details specific to each strategy are reported in the code and will be publicly available upon acceptance.
@@ 3.2 Robustness of classifiers to domain-shift
Tab. 3 shows the performance of classifiers trained using features extracted from each FM for diagnosis and breast density classification. For both tasks, Unified
classifiers trained on features from MammoCLIP outperform those based on GLORIA (, average relative improvement of +15.3%), CLIP (, +9.7%), and DinoV2 (, +13.3%), highlighting the advantages of pre-training on modality-specific data compared to domain-related (i.e. X-rays) or natural images. It is worth mentioning that MedCLIP-based classifiers exhibit notably poor performance across all tasks, suggesting that the extracted features may be predominantly noisy. Overall, classifiers have high performance when tested on the same dataset they were trained on (Indiv. (internal)), with average F1 scores of 0.73 and 0.53 for MammoCLIP on diagnosis and breast density, respectively. However, when tested on other datasets (Indiv. (external)), a drastic drop in F1 scores (up to for diagnosis) was remarkable, potentially due to overfitting of dataset-specific characteristics encoded in MammoCLIP’s features.
@@ 3.3 Effectiveness of bias and domain-adaptation strategies
Unified
classifiers show performance similar to individual classifiers on their test sets (internal), with F1 score variations of ±15%. Aggregating datasets effectively improves generalization compared to individual classifiers (overall F1 score improved by and , with Unified
vs. Indiv.). However, Unified
classifiers exhibit performance disparities, with F1 score standard deviations of 0.15 across test datasets, suggesting that such aggregation cannot fully mitigate biases. DANN
shows slightly lower overall F1 scores than Unified
classifiers, especially for the diagnosis task. While this technique aims to learn domain-invariant representations, it seems to do so at the cost of overall performance. Similar observations can be made for FairDisCO
, FADES
, and MOE
. For Breast density, GroupDRO
produces consistently tighter F1 score distributions than Unified
classifiers with standard deviations of 0.08, 0.06, and 0.08 for MammoCLIP, CLIP, and DinoV2, respectively, indicating reductions of disparities across test datasets. Additionally, it outperforms the Unified
classifier across FMs, e.g. MammoCLIP with a relative improvement of 12% (). For diagnosis, mitigation strategies do not improve overall performance nor reduce disparities compared to Unified
, likely due to variability in diagnostic label availability across datasets, e.g. no benign samples for MMD.
@@ 3.4 Bias and domain-adaptation in under-represented subgroups
Certain subgroups are under-represented in our datasets, e.g. breast density classes A and D representing 11 and 9% of the dataset and age <40 (10%) and >70 (10%), and are unequally represented across datasets (see Tab. 2). Fig.2 (top) illustrates prediction disparities across datasets and breast density classes, where indicates fair performance. GroupDRO
and DANN
achieve the most fair performance for all breast density classes (, and across breast density classes and datasets), aligning with DANN
’s domain-invariant feature learning strategy. However, for DANN
, this fairness comes at the cost of performance (see Tab. 3). GroupDRO
stabilizes performance across breast density classes, notably improving prediction for class A, which other classifiers struggled with. This ability to learn across domains, while favoring under-represented subgroups, is critical for extreme breast densities (A, D) and age groups due to their strong interplay [8]. Fig. 2 (bottom) presents samples from different subgroups and classifiers’ successes and failures in breast density classification. Variations in contrast, texture, and patterns across classes and datasets may introduce spurious correlations, underscoring the need for fairness-aware strategies. GroupDRO
seems an effective strategy to mitigate these biases and could be further refined to incorporate more fine-grained attributes.